# The Cramér-Rao approach and global quantum estimation of bosonic states

### v4

@@ V.3.1 Probe state as a binomial distribution in the number basis
$\frac{1}{n+2}\sum_{m=-j}^{j}\frac{j+m+1}{2j+1}{n\choose j+m}p^{j+m}(1-p)^{j-m}$ (49)
@@
$\frac{1}{n}\sum_{m=-j}^{j}\frac{j-m}{2j+1}{n\choose j+m}p^{j+m}(1-p)^{j-m}$ (51)
@@
V.3.2 Probe state as a geometric distribution in the number basis
@@
$\frac{1}{n+2}\sum_{m=-j}^{j}\frac{j+m+1}{2j+1}\frac{r-1}{r^{2j+1}-1}r^{j+m}$ (53)
@@
When $r>1$ and $n$ goes to infinity, LHS of (47) approaches $1$ and and RHS of (47) approaches 00. With sufficiently large $n$, the condition (47) holds. Then, the maximum of $R(M)$ is $\frac{n+1}{n+2}\Big{(}\frac{1}{n+1}+\frac{1}{n+1}(\frac{nr^{n+1}+1}{r^{n+1}-1}-\frac{1}{r-1})\Big{)}$, which converges to $1$, where $R(M)$ is defined in (43). When $M_{n}$ is the optimal estimator, we show in Appendix E that the corresponding error function is
@@
V.3.3 Probe state as a delta distribution in the number basis
@@ VI Discussion
In the case of unitary estimation of a single parameter, it is known that the optimal Cramér-Rao type bound cannot be attained with global estimation [13, 12, 60]. The papers [13, 12, 60] considers the optimization of the initial state for the estimation of the unitary. However, this paper considers the state estimation with a fixed initial state. Furthermore, the paper [12] showed that the optimal Cramér-Rao type bound cannot be attained even under the problem of local minimax estimation even under the setting of asymptotically many probe states used for global estimation of unitary channels. This phenomenon relates to the Heisenberg scaling under the unitary estimation. In unitary estimation, while the optimal Cramér-Rao type bound has the same order as the error function for optimum minimax estimation [61, 62, 13, 12, 60], they differ in the coefficients of their leading order terms. Moreover, it was shown that the input state of the optimal Cramér-Rao type bound and the optimum minimax estimation are different. For example, although the noon state realizes the optimal Cramér-Rao type bound, it does not work for global estimation [63, Section VI]. Our results add to the literature of examples where the behavior of Cramér-Rao type bounds differs from the optimum minimax estimation, particularly with regards to the quantum estimation of bosonic states both in a single-parameter and a multi-parameter setting.
@@ Acknowledgement
MH is supported in part by the National Natural Science Foundation of China (Grant No. 62171212). YO acknowledges support from EPSRC (Grant No. EP/W028115/1).
Appendix A Local estimation under the general unitary model
To show several relations in Section V, we consider the spin $j$ system ${\cal H}_{j}$ spanned by $\{|j;m\rangle\}_{m=-j}^{j}$. For simplicity, if $j$ is clear from the context, we denote $|j;m\rangle$ as $|m\rangle$.
@@ Appendix A Local estimation under the general unitary model
$J_{+}:=$ $\sum_{m=-j}^{j-1}\sqrt{j(j+1)-m(m+1)}|m+1\rangle\langle m|$ (59)
@@
$J_{-}:=$ $\sum_{m=-j}^{j-1}\sqrt{j(j+1)-m(m+1)}|m\rangle\langle m+1|$ (61)
@@
$i[\rho,J_{1}]=\frac{i}{2}[\rho,(J_{+}+J_{-})]=\frac{i}{2}(-J_{+}\rho+\rho J_{+}-J_{-}\rho+\rho J_{-})$ (69)
$i[\rho,J_{2}]=\frac{1}{2}[\rho,(J_{+}-J_{-})]=\frac{1}{2}(-J_{+}\rho+\rho J_{+}+J_{-}\rho-\rho J_{-}).$ (70)
@@
$K_{+}:=$ $\sum_{m=-j}^{j-1}\frac{2(p_{m+1}-p_{m})}{p_{m+1}+p_{m}}\sqrt{(j-m)(j+m+1)}|m+1\rangle\langle m|$ (71)
$K_{-}:=$ $\sum_{m=-j}^{j-1}\frac{2(p_{m+1}-p_{m})}{p_{m+1}+p_{m}}\sqrt{(j-m)(j+m+1)}|m\rangle\langle m+1|.$ (72)
@@
$F_{1,1}=\mathrm{Tr}\rho L_{1}^{2}$ (76)
$=$ $\sum_{m=-j}^{j-1}\Big{(}\frac{4(p_{m+1}-p_{m})^{2}}{(p_{m+1}+p_{m})^{2}}(j-m)(j+m+1)p_{m}$ (77)
$+\frac{4(p_{m+1}-p_{m})^{2}}{(p_{m+1}+p_{m})^{2}}(j-m)(j+m+1)p_{m+1}\Big{)}$ (78)
$=$ $\sum_{m=-j}^{j-1}\frac{4(p_{m+1}-p_{m})^{2}}{p_{m+1}+p_{m}}(j-m)(j+m+1)$ (79)
$F_{2,2}=\mathrm{Tr}\rho L_{2}^{2}$ (80)
$=$ $\sum_{m=-j}^{j-1}\frac{4(p_{m+1}-p_{m})^{2}}{p_{m+1}+p_{m}}(j-m)(j+m+1)$ (81)
@@
Appendix B Local estimation with the binomial distribution: Proof of (38)
@@
$\langle j+\frac{1}{2};m+\frac{1}{2}|\frac{1}{2},\frac{1}{2};j,m\rangle^{2}=\frac{(2j+2)(2j)!(j+m)!(j-m)!(j+m+1)!(j-m)!}{(2j+2)!((j+m)!(j-m)!)^{2}}$ (86)
@@
$F_{1,1}=$ $\sum_{m=-j}^{j-1}\frac{4(p_{m+1}-p_{m})^{2}}{p_{m+1}+p_{m}}(j-m)(j+m+1)$ (88)
$=$ $\sum_{m=-j}^{j-1}\frac{4({n\choose j+m+1}p^{j+m+1}(1-p)^{j-m-1}-{n\choose j+m}p^{j+m}(1-p)^{j-m})^{2}}{{n\choose j+m+1}p^{j+m+1}(1-p)^{j-m-1}+{n\choose j+m}p^{j+m}(1-p)^{j-m}}$ (89)
$\cdot(j-m)(j+m+1)$ (90)
$=$ $\sum_{m=-j}^{j-1}\frac{4(\frac{j-m}{j+m+1}\frac{p}{1-p}-1)^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m}}{\frac{j-m}{j+m+1}\frac{p}{1-p}+1}(j-m)(j+m+1)$ (91)
$=$ $\sum_{m=-j}^{j-1}\frac{4((j-m)\frac{p}{1-p}+(j+m+1))^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m}}{(j-m)\frac{p}{1-p}+(j+m+1)}(j-m)$ (92)
$=$ $\sum_{m=-j}^{j-1}\frac{4((j-m)p-(j+m+1)(1-p))^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m-1}}{(j-m)p+(j+m+1)(1-p)}(j-m)$ (93)
$=$ $\sum_{m=-j}^{j-1}\frac{4(-(j+m+1)+(2j+1)p)^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m-1}}{j+m+1-(2m+1)p}(j-m)$ (94)
$=$ $\sum_{k=0}^{n-1}\frac{4(-(k+1)+(n+1)p)^{2}{n\choose k}p^{k}(1-p)^{2n-k-1}}{k+1-(2k-n+1)p}(n-k)$ (95)
@@
$F_{1,1}=$ $n^{2}\mathbb{E}_{0,p(1-p)}\Big{[}\frac{4(-p-\frac{X}{\sqrt{n}}+(1+\frac{1}{n})p)^{2}(1-p)^{-1}(1-p-\frac{X}{\sqrt{n}})}{p+\frac{X}{\sqrt{n}}+\frac{1}{n}-(2p+2\frac{X}{\sqrt{n}}-1+\frac{1}{n})p}\Big{]}$ (98)
$=$ $n^{2}\mathbb{E}_{0,p(1-p)}\Big{[}\frac{4(\frac{X}{\sqrt{n}}-\frac{p}{n})^{2}(1-p)^{-1}(1-p-\frac{X}{\sqrt{n}})}{-2p^{2}+2p+(1-2p)\frac{X}{\sqrt{n}}+\frac{1-p}{n}}\Big{]}$ (99)
@@
Appendix C Local estimation with the geometric distribution
@@
$-\rho\circ\frac{2(r+1)}{r-1}J_{2}=$ $i[\rho,J_{1}]$ (107)
$\rho\circ\frac{2(r+1)}{r-1}J_{1}=$ $i[\rho,J_{2}].$ (108)
@@
$C^{\rm R}_{\bm{\theta}}[I]=\mathrm{Tr}F^{-1}+\frac{1}{2}\mathrm{Tr}|F^{-1}DF^{-1}|$ (112)
$\cong$ $\frac{2(r+1)}{n(r-1)}+\frac{2}{n}=\frac{2(r+1)+2(r-1)}{n(r-1)}=\frac{4r}{n(r-1)},$ (113)
@@
Appendix D Proof of Theorem 1
For reader’s convenience, we give a proof for Theorem 1 as a special case of [52, Theorem 4.6].

### v5

@@ 5.3.1 Probe state as a binomial distribution in the number basis
$\frac{1}{n+2}\sum_{m=-j}^{j}\frac{j+m+1}{2j+1}{n\choose j+m}p^{j+m}(1-p)^{j-m}$
@@
$\frac{1}{n}\sum_{m=-j}^{j}\frac{j-m}{2j+1}{n\choose j+m}p^{j+m}(1-p)^{j-m}$
@@
5.3.2 Probe state as a geometric distribution in the number basis
@@
$\frac{1}{n+2}\sum_{m=-j}^{j}\frac{j+m+1}{2j+1}\frac{r-1}{r^{2j+1}-1}r^{j+m}$
@@
When $r>1$ and $n$ goes to infinity, LHS of (47) approaches $1$ and and RHS of (47) approaches 00. With sufficiently large $n$, the condition (47) holds. Then, the maximum of $R(M)$ is $\frac{n+1}{n+2}\Big{(}\frac{1}{n+1}+\frac{1}{n+1}(\frac{nr^{n+1}+1}{r^{n+1}-1}-\frac{1}{r-1})\Big{)}$, which converges to $1$, where $R(M)$ is defined in (43). When $M_{n}$ is the optimal estimator, we show in Appendix F that the corresponding error function is
@@
5.3.3 Probe state as a delta distribution in the number basis
@@ 6 Discussion
In the case of unitary estimation of a single parameter, it is known that the optimal Cramér-Rao type bound cannot be attained with global estimation [13, 12, 72]. The papers [13, 12, 72] consider the optimization of the initial state for the estimation of the unitary. However, this paper considers the state estimation with a fixed initial state. Furthermore, the paper [12] showed that the optimal Cramér-Rao type bound cannot be attained even under the problem of local minimax estimation even under the setting of asymptotically many probe states used for global estimation of unitary channels. This phenomenon relates to the Heisenberg scaling under the unitary estimation. In unitary estimation, while the optimal Cramér-Rao type bound has the same order as the error function for optimum minimax estimation [73, 74, 13, 12, 72], they differ in the coefficients of their leading order terms. Moreover, it was shown that the input state of the optimal Cramér-Rao type bound and the optimum minimax estimation are different. For example, although the noon state realizes the optimal Cramér-Rao type bound, it does not work for global estimation [75, Section VI]. Our results add to the literature of examples where the behavior of Cramér-Rao type bounds differs from the optimum minimax estimation, particularly with regards to the quantum estimation of bosonic states both in a single-parameter and a multi-parameter setting.
@@
Next, we discuss the global attainability of the Cramér-Rao bound, which represents the minimum of the mean squared error (MSE) sum under the local unbiasedness condition. In the single-parameter case, the Cramér-Rao bound coincides with the optimal error under the local unbiasedness condition if and only if the state family is an exponential family in the sense of SLD (symmetric logarithmic derivative) [17, 18][16, Theorem 6.7]. In the multiple-parameter case, it is difficult to obtain the same assertion due to the complexity of problems associated with non-commutativity in minimizing the MSE sum under the local unbiasedness condition. However, its connection with exponential families has been discussed in recent literature [44]. At least in the classical case, under certain regularity conditions on the distribution family, being an exponential family is equivalent to the Cramér-Rao bound being globally attainable.
It is conceivable that as the state family approaches an exponential family, the Cramér-Rao bound becomes globally attainable. The quantum Gaussian states family is shown to be a non-commutative exponential family [44, Section 6]. For instance, in state estimation, the Cramér-Rao bound is globally attainable when the number of copies is large. As shown in [76, 77, 4], in this case, it asymptotically approaches the quantum Gaussian states family, albeit locally. Furthermore, the state family treated in Subsection 5.3.2 also asymptotically approaches the quantum Gaussian states family, as shown in [5, Theorem 7] in the limit of large $j$. Therefore, as confirmed in (53), the Cramér-Rao bound is asymptotically globally attainable. In other words, it is safer to assume that the Cramér-Rao bound is not globally attainable unless the situation approaches an exponential family.
Finally, we discuss the relation between the unbiased condition and the group covariant approach discussed in Section 5.3. Since our state space forms a compact set, any coordinate cannot cover the whole of the parameter space smoothly. Now, we employ the parameter space $\{(\theta_{1},\theta_{2})||\theta|\leq\pi\}$, and this space has discontinuity at the state $|\frac{1}{2};-\frac{1}{2}\rangle\langle\frac{1}{2};-\frac{1}{2}|$. Due to this issue, any covariant estimator including our optimal estimator has bias except for the origin $\theta=0$ under this parametrization because the center of the measurement outcome is shifted to the direction of the origin. Since the Cramér-Rao approach employs only the locally unbiased condition and Section 5.1 discusses the error at the origin, it is more important whether our optimal estimator satisfies the locally unbiased condition. However, it is difficult to check whether a covariant estimator satisfies the locally unbiased condition because the quantities appearing in this condition cannot be easily handled in the group covariant framework. Nevertheless, the state family treated in Subsection 5.3.2 also asymptotically approaches the quantum Gaussian states family in the limit of large $j$, as mentioned the above. Since the optimal covariant estimator for the quantum Gaussian states family satisfies the unbiasedness condition [2, Chapter 6], the optimal covariant estimator asymptotically satisfies the unbiasedness condition in this case.
@@ Acknowledgement
MH is supported in part by the National Natural Science Foundation of China (Grant No. 62171212), and the General R&D Projects of 1+1+1 CUHK-CUHK(SZ)-GDST Joint Collaboration Fund (Grant No. GRDP2025-022). YO acknowledges support from EPSRC (Grant No. EP/W028115/1).
Appendix A Derivation of (41)
We have
$U_{\theta}:=\exp(i(\theta_{1}J_{1}+\theta_{2}J_{2}))=\exp\left(\begin{pmatrix}0&\frac{i\theta_{1}+\theta_{2}}{2}\\ \frac{i\theta_{1}-\theta_{2}}{2}&0\end{pmatrix}\right)$
$=$ $\exp\left(\left(\begin{array}[]{cc}0&ie^{-i\phi}\frac{|\theta|}{2}\\ ie^{i\phi}\frac{|\theta|}{2}&0\end{array}\right)\right)$ (58)
$=$ $\exp\left(\left(\begin{array}[]{cc}1&0\\ 0&-ie^{i\phi}\end{array}\right)\left(\begin{array}[]{cc}0&\frac{|\theta|}{2}\\ -\frac{|\theta|}{2}&0\end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&ie^{-i\phi}\end{array}\right)\right)$ (65)
$=$ $\left(\begin{array}[]{cc}1&0\\ 0&-ie^{i\phi}\end{array}\right)\exp\left(\left(\begin{array}[]{cc}0&|\theta|\\ -|\theta|&0\end{array}\right)\right)\left(\begin{array}[]{cc}1&0\\ 0&ie^{-i\phi}\end{array}\right)$ (72)
$=$ $\left(\begin{array}[]{cc}1&0\\ 0&-ie^{i\phi}\end{array}\right)\left(\begin{array}[]{cc}\cos\frac{|\theta|}{2}&\sin\frac{|\theta|}{2}\\ -\sin\frac{|\theta|}{2}&\cos\frac{|\theta|}{2}\end{array}\right)\left(\begin{array}[]{cc}1&0\\ 0&ie^{-i\phi}\end{array}\right)$ (79)
$=$ $\left(\begin{array}[]{cc}\cos\frac{|\theta|}{2}&ie^{-i\phi}\sin\frac{|\theta|}{2}\\ ie^{i\phi}\sin\frac{|\theta|}{2}&\cos\frac{|\theta|}{2}\end{array}\right).$ (82)
Thus,
$U_{\theta}|\frac{1}{2};\frac{1}{2}\rangle=\cos\frac{|\theta|}{2}|\frac{1}{2};\frac{1}{2}\rangle+ie^{i\phi}\sin\frac{|\theta|}{2}|\frac{1}{2};-\frac{1}{2}\rangle.$ (83)
Hence,
$R(\theta,\hat{\theta})=$ $|\langle\frac{1}{2};\frac{1}{2}|U_{\theta}^{\dagger}U_{\hat{\theta}}|\frac{1}{2};\frac{1}{2}\rangle|^{2}$
$=$ $\Big{|}\cos\frac{|\theta|}{2}\cos\frac{|\hat{\theta}|}{2}+e^{i(-\phi+\hat{\phi})}\sin\frac{|\theta|}{2}\sin\frac{|\hat{\theta}|}{2}\Big{|}^{2}.$ (84)
Appendix B Local estimation under the general unitary model
To show several relations in Section 5, we consider the spin $j$ system ${\cal H}_{j}$ spanned by $\{|j;m\rangle\}_{m=-j}^{j}$. For simplicity, if $j$ is clear from the context, we denote $|j;m\rangle$ as $|m\rangle$.
@@ Appendix B Local estimation under the general unitary model
$J_{+}:=$ $\sum_{m=-j}^{j-1}\sqrt{j(j+1)-m(m+1)}|m+1\rangle\langle m|$
@@
$J_{-}:=$ $\sum_{m=-j}^{j-1}\sqrt{j(j+1)-m(m+1)}|m\rangle\langle m+1|$
@@
$i[\rho,J_{1}]$ $=\frac{i}{2}[\rho,(J_{+}+J_{-})]$
$=\frac{i}{2}(-J_{+}\rho+\rho J_{+}-J_{-}\rho+\rho J_{-})$ (92)
$i[\rho,J_{2}]$ $=\frac{1}{2}[\rho,(J_{+}-J_{-})]$
$=\frac{1}{2}(-J_{+}\rho+\rho J_{+}+J_{-}\rho-\rho J_{-}).$ (93)
@@
$K_{+}:=$ $\sum_{m=-j}^{j-1}\frac{2(p_{m+1}-p_{m})}{p_{m+1}+p_{m}}$
×$\quad\quad\times\sqrt{(j-m)(j+m+1)}|m+1\rangle\langle m|$ (94)
$K_{-}:=$ $\sum_{m=-j}^{j-1}\frac{2(p_{m+1}-p_{m})}{p_{m+1}+p_{m}}$
×$\quad\quad\times\sqrt{(j-m)(j+m+1)}|m\rangle\langle m+1|.$ (95)
@@
$F_{1,1}=\mathrm{Tr}\rho L_{1}^{2}$
$=$ $\sum_{m=-j}^{j-1}\Big{(}\frac{4(p_{m+1}-p_{m})^{2}}{(p_{m+1}+p_{m})^{2}}(j-m)(j+m+1)p_{m}$
$+\frac{4(p_{m+1}-p_{m})^{2}}{(p_{m+1}+p_{m})^{2}}(j-m)(j+m+1)p_{m+1}\Big{)}$
$=$ $\sum_{m=-j}^{j-1}\frac{4(p_{m+1}-p_{m})^{2}}{p_{m+1}+p_{m}}(j-m)(j+m+1),$ (99)
$F_{2,2}=\mathrm{Tr}\rho L_{2}^{2}$
$=$ $\sum_{m=-j}^{j-1}\frac{4(p_{m+1}-p_{m})^{2}}{p_{m+1}+p_{m}}(j-m)(j+m+1),$ (100)
@@
Appendix C Local estimation with the binomial distribution: Proof of (37)
@@
$\langle j+\frac{1}{2};m+\frac{1}{2}|\frac{1}{2},\frac{1}{2};j,m\rangle^{2}=\frac{(2j+2)(2j)!(j+m)!(j-m)!(j+m+1)!(j-m)!}{(2j+2)!((j+m)!(j-m)!)^{2}}$
@@
$F_{1,1}=$ $\sum_{m=-j}^{j-1}\frac{4(p_{m+1}-p_{m})^{2}}{p_{m+1}+p_{m}}(j-m)(j+m+1)$
$=$ $\sum_{m=-j}^{j-1}\frac{4({n\choose j+m+1}p^{j+m+1}(1-p)^{j-m-1}-{n\choose j+m}p^{j+m}(1-p)^{j-m})^{2}}{{n\choose j+m+1}p^{j+m+1}(1-p)^{j-m-1}+{n\choose j+m}p^{j+m}(1-p)^{j-m}}$
$\cdot(j-m)(j+m+1)$
$=$ $\sum_{m=-j}^{j-1}\frac{4(\frac{j-m}{j+m+1}\frac{p}{1-p}-1)^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m}}{\frac{j-m}{j+m+1}\frac{p}{1-p}+1}(j-m)(j+m+1)$
$=$ $\sum_{m=-j}^{j-1}\frac{4((j-m)\frac{p}{1-p}+(j+m+1))^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m}}{(j-m)\frac{p}{1-p}+(j+m+1)}(j-m)$
$=$ $\sum_{m=-j}^{j-1}\frac{4((j-m)p-(j+m+1)(1-p))^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m-1}}{(j-m)p+(j+m+1)(1-p)}(j-m)$
$=$ $\sum_{m=-j}^{j-1}\frac{4(-(j+m+1)+(2j+1)p)^{2}{n\choose j+m}p^{j+m}(1-p)^{j-m-1}}{j+m+1-(2m+1)p}(j-m)$
$=$ $\sum_{k=0}^{n-1}\frac{4(-(k+1)+(n+1)p)^{2}{n\choose k}p^{k}(1-p)^{2n-k-1}}{k+1-(2k-n+1)p}(n-k)$
@@
$F_{1,1}=$ $n^{2}\mathbb{E}_{0,p(1-p)}\Big{[}\frac{4(-p-\frac{X}{\sqrt{n}}+(1+\frac{1}{n})p)^{2}(1-p)^{-1}(1-p-\frac{X}{\sqrt{n}})}{p+\frac{X}{\sqrt{n}}+\frac{1}{n}-(2p+2\frac{X}{\sqrt{n}}-1+\frac{1}{n})p}\Big{]}$
$=$ $n^{2}\mathbb{E}_{0,p(1-p)}\Big{[}\frac{4(\frac{X}{\sqrt{n}}-\frac{p}{n})^{2}(1-p)^{-1}(1-p-\frac{X}{\sqrt{n}})}{-2p^{2}+2p+(1-2p)\frac{X}{\sqrt{n}}+\frac{1-p}{n}}\Big{]}$
@@
Appendix D Local estimation with the geometric distribution
@@
$-\rho\circ\frac{2(r+1)}{r-1}J_{2}=i[\rho,J_{1}],\quad\rho\circ\frac{2(r+1)}{r-1}J_{1}=i[\rho,J_{2}].$ (114)
@@
$C^{\rm R}_{\bm{\theta}}[I]=\mathrm{Tr}F^{-1}+\frac{1}{2}\mathrm{Tr}|F^{-1}DF^{-1}|\cong\frac{2(r+1)}{n(r-1)}+\frac{2}{n}=\frac{2(r+1)+2(r-1)}{n(r-1)}=\frac{4r}{n(r-1)},$ (118)
@@
Appendix E Proof of Theorem 1
For reader’s convenience, we give a proof for Theorem 1 as a special case of [64, Theorem 4.6].